OpenStudy (anonymous):

How to find period of full oscillation of H-Br bond ? A spectrum of vibrational rotation spectrum is given in units of energy, eV.

OpenStudy (irishboy123):

the energy should be quantised according to E = (n + 1/2) h f where h = Planck, f = freq you can use this equation and the energy bands to calc f and T = 1/f.